OpenStudy (anonymous):

An equation is shown below: 6(2x - 11) + 15 = 21 Write the steps you will use to solve the equation and explain each step.

OpenStudy (anonymous):

(Essay Worth 5 points

OpenStudy (marihelenh):

1. You would use the distributive property to get rid of parenthesis. 2. You would add like terms on the left side. 3. Subtract or add whatever is on the left side that does not have the x term. 4. Divide both sides by the coefficient in front of x.

OpenStudy (marihelenh):

\[6x-66+15=21\] \[6x-51=21\] \[6x=72\] \[x=12\]

imqwerty (imqwerty):

1 Subtract 15 from both sides 6(2x−11)=21−15 2 Simplify 21−15 to 6 6(2x−11)=6 3 Divide both sides by 6 2x−11=1 4 Add 11 to both sides 2x=1+11 5 Simplify 1+11 to 12 2x=12 6 Divide both sides by 2 x=12/2 7 Simplify 12/2 to 6 x=6